Q2. Crown Ltd has the following book value capital structure. Equity capital (shares of Rs 10 par value each) Rs 15 crore, 12% Preference capital (Rs 100 par value each) Rs 1 crore. Retained earnings Rs 20 crore, 11.5% Debentures (Rs 100 par value each) 10 crore and 11% Term loan Rs 12.5 crore. The next year expected dividend on equity is Rs 3.6 per share and has an expected growth rate of 7%. The market value is Rs 40/share. Preference stock, redeemable after 10 years is currently trading at Rs 75 per share. Debentures, trading at Rs 80 are redeemable after 6 years. Corporate tax rate is 40%. Calculate the WACC as per book value weights. Comment on the relevance of calculation of WACC.

1
Expert's answer
2021-05-17T12:12:12-0400

WACC as per book value weights


Book value of Equity & Retained Earnings = Rs.15 crores+ Rs.20 crores =Rs.35 crores

Book Value of preference Shares = Rs. 1 crores

Book Value of Debentures =Rs. 10 crores

Book value of Term loan = Rs.12.5 crores


Cost of Equity = Next year Dividend/Share price + growth rate =3.6/40+0.07 = 0.1600 or 16.00%


Cost of Preference Shares = Dividend/ market price of Preference Share = Rs.100*12%/Rs.75 = 0.1600 or 16.00%


Pretax cost of Debentures(r) is given by (assuming annual coupon)


11.5/r*(1-1/(1+r)^6)+100/(1+r)^6 =80


Solving r =0.1708 or 17.08%


Using Book value weights


WACC = "\dfrac{35}{(35+1+10+12.5)}*\dfrac{16}{100}+\dfrac{1}{(35+1+10+12.5)}*\dfrac{16}{100}+\dfrac{10}{(35+1+10+12.5)}*\dfrac{17.08}{100}*(1-0.4) +\dfrac{12.5}{(35+1+10+12.5)}*\dfrac{11}{100}*(1-0.4)"

=13.008% or 13.01%


WACC as per market value weights


Market value of Equity & Retained Earnings = No of shares* Market price per share + Rs.20 crores


=15 crores/10 *40 + Rs.20 crores = Rs.80 crores


Market Value of preference Shares = No of preference shares* Market price per preference share


=Rs1 crores /100 * 75 = Rs.0.75 crores


Market Value of Debentures =No of Debentures* Market price per Debenture


=Rs10 crores /100 * 80 = Rs.8 crores


Market value of Term loan = Rs.12.5 crores


Using Market value weights


"WACC = \dfrac{80}{(80+0.75+8+12.5)}*\dfrac{16}{100}+\dfrac{0.75}{(80+0.75+8+12.5)}*\dfrac{16}{100}+\dfrac{8}{(80+0.75+8+12.5)}*\dfrac{17.08}{100}*(1-0.4)+\dfrac{12.5}{(80+0.75+8+12.5)}*\dfrac{11}{100}*(1-0.4)"


=14.3852% or 14.39%